OpenStudy (brucebaner):

Tickets to the museum cost $18 for adults and $12 for children. A group of 20 people went to the museum, and the tickets cost $324. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y=20 18x+12y=324 How many adults and how many children were in the group? A. 8 adults and 12 children B. 10 adults and 10 children C. 12 adults and 8 children D. 14 adults and 6 children

OpenStudy (zale101):

\(\Large x+y=20\) solve for y then plug the equation for y into this equation \(\Large 18x+12\color{red}y=324\)

OpenStudy (zale101):

Clear? :)

OpenStudy (zale101):

i'm solving the systems of equation by using substitution

OpenStudy (brucebaner):

okay

OpenStudy (zale101):

to know how many adults \(\color {red}y\) and how many children \(\color{blue} x\) were in the group of 20 people, we must combine the two equations by using the systems of equation as the direction states. I chose doing substitution because it's more understanding. You know how to solve by substitution?

OpenStudy (zale101):

\(\large Step:1\) \(\color{blue} x+\color{red}y=20\) solve for \(\color{red}y\)

OpenStudy (brucebaner):

okay

OpenStudy (brucebaner):

so how do i do what plz help me

OpenStudy (zale101):

\(\Large \color{blue}x+\color{red}y=20\) subtract x to both sides \(\Large \color{blue}{x-x}+\color{red}{y}=20-\color{blue}{x}\) you'll end up with \(\Large\color{red}y=20-\color{blue}x\)

OpenStudy (zale101):

plug \(\color{red} y\) which is \(20-\color{blue}{x}\) to y in the equation \(\large 18x+12\color{red}y=324\)

OpenStudy (zale101):

\(\Large 18x+12(20-\color{blue}x)=324\) clear?

OpenStudy (brucebaner):

yes

OpenStudy (zale101):

Can you solve for x? Like, distributing and combining like terms? \(\Large 18x+\color{red}{12(20-x)}=324\) distribute the red shaded color, you know how to do that?

OpenStudy (brucebaner):

no

OpenStudy (zale101):

solve for x in this case, x represents how many adults went to the museum

OpenStudy (zale101):

 \(\Large 18x+240-12x=324\)